Find the sum of the exponential series 96 + 24 + 6 + ....
Answer by sachi(548)

the infinite(correct word) series 96 + 24 + 6 + ....
here the 1st term=t1=96
& 2nd term t2=24
so common ratio r=24/96=1/4=0.25 <1
so the sum =a/(r-1)=96/(1-0.25)=96/0.75=128